John is presently 12 years older than Fred. Four years ago John was twice as old as Fred. How old is each now?
Answer by macston(5194)

J=John's age=F+12 years: F=Fred's age
J-4 years=2(F-4 years) Substitute for J
F+12 years-4 years=2(F-4 years)
F+8=2F-8 years Add 8 years to each side.
F+16 years=2F Subtract F from each side.
16 years=F ANSWER: Fred is currently 16 years old.
J=F+12 years=16 years+12 years=28 years ANSWER: John is currently 28 years old.
CHECK:
J-4 years=2(F-4 years)
28 years-4 years=2(16 years-4 years)
24 years=2(12 years)
24 years=24 years